OpenStudy (anonymous):

What is the greatest whole number that must be a divisor of the product of any three consecutive positive integers?

DivineSolar (divinesolar):

This deals with integers correct?

OpenStudy (anonymous):

Yes and number theory to be precise

DivineSolar (divinesolar):

Let me see what i can do...

DivineSolar (divinesolar):

Hmm i am afraid my knowledge on this is to short... I am sorry @danyboy9169 I dont have much knowledge with gwh yet..

DivineSolar (divinesolar):

GWH=Greatest whole number.

OpenStudy (anonymous):

Yes..I think the ans is 6 I got it right,, it is easy to do ..

DivineSolar (divinesolar):

Awesome nice to know you got it correct!

OpenStudy (anonymous):

see if you take any three numbers and multiply them they are divisible by 6 which is the GWH.... I just did trial and error and got it right..thinking too deeply puts me on the wrong foot

DivineSolar (divinesolar):

Cool :)! I need to go help others now :)!

OpenStudy (thomas5267):

How to prove the lower bound is 6?

OpenStudy (amistre64):

n(n-1)(n+1) n(n^2-1) n^3-n seems to be all i got at the moment :)

OpenStudy (thomas5267):

The upper bound is 6 because 1*2*3=6.

OpenStudy (anonymous):

1*2*3= 6 4*5*6= 120 6*7*8=336 Try any 3 consecutive #s they are always divisible by 6 We know that at least one of our three consecutive positive integers must be a multiple of $2$ since every other integer in a list of consecutive integers is divisible by $2$. Similarly, one of our three consecutive integers must also be divisible by $3$. Thus, the product of our three integers must be divisible by $2 \cdot 3 = 6$. By choosing the example where our three consecutive integers are $1$, $2$, and $3$ and their product is $6$, we see that $\boxed{6}$ is indeed the greatest whole number that must be a factor of the product of any three consecutive positive integers.

ganeshie8 (ganeshie8):

**\[n^{\phi(6)} \equiv 1 \pmod{6} \implies n^2\equiv 1 \pmod{6} \implies n^3\equiv n \pmod{6}\]

OpenStudy (thomas5267):

n(n+1)(n+2) -> Three numbers, one of them will be 0 mod 3. n(n+1)(n+2) -> More than two numbers, at least one of them will be 0 mod 2. Therefore, n(n+1)(n+2)=0 (mod 6)

OpenStudy (thomas5267):

The upper bound and lower bound are both 6. Therefore, the number is 6.

OpenStudy (anonymous):

wow! great.. ganeshie8 and thanks Thomas5267, now it is more clear

OpenStudy (thomas5267):

Suddenly realised that the middle number between the twin primes must be divisible by 6.